OpenStudy (anonymous):

Solve each equation. Write all proposed solutions. Cross out those that are extraneous.

OpenStudy (anonymous):

\[\sqrt{2m+4}-\sqrt{m+3}=1\]

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

start with \[\sqrt{2m+4}=1+\sqrt{m+3}\] then square both sides

OpenStudy (anonymous):

My thoughts exactly!

OpenStudy (anonymous):

you get \[2m+4=1+m+3+2\sqrt{m+3}\] now get the radical by itself on the right and square again

OpenStudy (anonymous):

\[2m+4=m+4+2\sqrt{m+3}\] \[m=2\sqrt{m+3}\] not really so bad actually

OpenStudy (anonymous):

btw i hope it is clear what i did took \[(\sqrt{2m+4})^2=(1+\sqrt{m+3})^2\]

OpenStudy (anonymous):

I could follow it, just barely but i got it. Thank you again!